# PsycoLLM: Enhancing LLM for Psychological Understanding and Evaluation

### v1

@@ Abstract
Mental health has attracted substantial attention in recent years and LLM can be an effective technology for alleviating this problem owing to its capability in text understanding and dialogue. However, existing research in this domain often suffers from limitations, such as training on datasets lacking crucial prior knowledge and evidence, and the absence of comprehensive evaluation methods. In this paper, we propose a specialized psychological large language model (LLM), named PsycoLLM, trained on a proposed high-quality psychological dataset, including single-turn QA, multi-turn dialogues enriched with prior knowledge and knowledge-based QA. Additionally, to compare the performance of PsycoLLM with other LLMs, we develop a comprehensive psychological benchmark based on authoritative psychological counseling examinations in China, which includes assessments of professional ethics, theoretical proficiency, and case analysis. The experimental results on the benchmark illustrates the effectiveness of PsycoLLM, which demonstrates superior performance compared to other LLMs.
1 Introduction
@@
Large language models (LLM) have attracted tremendous attention and experienced promising advancements, such as ChatGPT, KimiChat111https://kimi.moonshot.cn/, PaLM Chowdhery et al. (2023) , BaiChuan Yang et al. (2023) and LLama Touvron et al. (2023). These models have demonstrated significant advancements across a variety of tasks, including question answering (QA), summarization Zhang et al. (2023a), and information extraction Wei et al. (2023). Moreover, large language models (LLMs) have been effectively applied to diverse domains, such as the medical and legal domains Xiong et al. (2023); Cui et al. (2023); Chen et al. (2023a). In recent years, the escalating pressures from personal, and occupational demands have emerged as significant challenges to the mental health status of individuals. Since LLMs are trained on massive data and contain huge number of parameters, many experiments have shown that these models have the strong ability to understand the natural language and provide resaonable feedback. Hence, leveraging LLM to aid individuals in addressing psychological issues may represent an efficacious strategy for mitigating the impact of adverse mental health conditions and making it easier for seekers to obtain timely help.
@@
Many studies have been proposed in this area. For example, Lai et al. (2023) combined pre-trained LLMs with real-world professional QA data, with the purpose of providing online consultation services. Chen et al. (2023b) and Qiu et al. (2023a) constructed a multi-turn empathetic conversation dataset and used a base LLM for fine-tuning. Further, EmoLLM (2024) follow a similar way and fine-tune the open base model on the mental health dataset. Although these existing efforts have contributed significantly to the advancement of the field, they still exhibit certain limitations. First, most existing psychological fine-tuning datasets are generated by even larger language model such as ChatGPT. However, the generation process lacks adequate prior knowledge and evidence, potentially leading to the creation of datasets that deviate from real-world psychological communication scenarios. Second, owing to the lack of a comprehensive psychological benchmark, the evaluation of most studies in this area often relies on expert assessment based on subjective judgment or overly depends on other large models, which may not fully encompass the subtle performance. For instance, Zhang et al. (2023c) utilized questions generated by GPT-4 to evaluate model performance.
@@
Therefore, in this paper, we propose a psychological LLM (PsycoLLM) trained on a high-quality dataset tailored to the domain of psychology, encompassing single-turn QA, multi-turn dialogues, and knowledge-based QA. Specifically, for single-turn QA, we collect question-answer pairs from open platforms. For multi-turn dialogues, we develop an innovative pipeline with generation, evidence judgment, and refinement for producing high-quality data. For knowledge-based QA, we extract after-school exercises from psychology textbooks and use LLMs enhanced with agents to extract QA pairs from psychology books. Additionally, to explore the performance of the proposed model and exiting LLMs, we introduce a benchmark based on authoritative psychological counseling examinations in China for comparative analysis. The benchmark consists of three distinct components: professional ethics, theoretical proficiency, and case analysis. Professional ethics and theoretical proficiency are assessed through multiple-choice questions (MCQs) to evaluate the mastery of psychological principles, while the case analysis component emphasizes practical utility by incorporating both MCQs and QA sections to explore the application of the model in real-world scenarios. The results reveal several insightful observations: I) Some LLMs have achieved an accuracy rate surpassing 60% in MCQs, showing the potential of using LLM to provide psychological assistance, with the proposed PsycoLLM demonstrating superior performance among these models. II) LLMs display greater proficiency in professional ethics compared to their mastery of psychological theory.
@@
2 PsycoLLM
@@
In this section, we present the details of dataset including, single-turn QA, multi-turn dialogue and knowledge-based QA, respectively. The overview of dataset are shown in 1. Besides, we give the training process of PsycoLLM.
2.1 Single-Turn QA Construction
@@
There are several publicly accessible websites committed to establishing a psychology platform and offering online solutions for individuals seeking psychological assistance, such as Yixinli, Zhihu and so on. The data from these platforms can be regarded as real-world inquiries and the solutions provided by professionals in response to these inquiries. On the original website, users typically provide descriptions along with their questions, including background information. Subsequently, professional or experienced individuals offer detailed solutions or advice in response to these questions. We collect over 267,000 pairs of data from websites. Furthermore, we conduct rigorous data cleaning procedures to enhance the quality and relevance of the dataset, including:
removal of irrelevant content (e.g., advertisements)
deletion of data with less than 100 characters
removal of repeated punctuation
removal of answers with fewer than 5 likes
removal of responses from lower-level counselors or individual answers
Note that the respondents’ level may serve as an indicator of their extensive experience in assisting help-seekers, which, in turn, can be indicative of their professional competence. This metric provides a basis for assessing the quality of the provided information. The number of likes functions as a measure of public endorsement of the response. A high number of likes suggests a broad consensus among users regarding the validity and reliability of the answer. After remove such data, we obtain over 155k pairs for single-turn QA.
@@
We report detailed statistics of the processed data drawn from different aspects after processing. We divide the processed data into 9 major topics and 127 subtopics, as illustrated by the distribution of major topics depicted in Figure 3. It is observed that the number of reports related to emotional issues and regulation, as well as interpersonal relationships and socializing, is higher than those in other categories, each comprising more than 20% of the total. Following these topics, family and marriage, and personal growth and development are the next most significant areas, each accounting for more than 10% of the reports. The visualization of the collected data is depicted through word clouds, as shown in Figure 2, which provides an intuitive depiction of the frequency and prominence of specific terms within the dataset. The analysis underscores the pervasive connection of the majority of these data to everyday psychological counseling, further emphasizing the relevance and reasonability of the collected data.
2.2 Multi-Turn Dialogue Construction
Based on the extracted data in section 2.1, we select a subset of the highest upvoted answers for each question to generate multi-turn dialogue data using KimiChat. This process aims to emulate the behavior of a psychological counselor who progressively asks probing questions to gain insight into the user’s psychological state and requirements, thereby facilitating more effective assistance. The multi-turn dialogue mainly includes two roles: the person seeking help and psychological professionals. We employ a three-step pipeline to generate high-quality multi-turn dialogue data. First, we use an appropriate prompt to guide the KimiChat to construct multi-turn dialogue data between pre-defined roles according to the selected pairs. The original pairs serve as prior knowledge to guide KimiChat in generating multi-turn conversations, thereby enhancing the fidelity of the generated data to real-world conversations. Second, we employ an additional prompt to assess whether the answers in the multi-turn dialogue dataset are derived from the original context. If the majority of responses in the multi-turn dialogue can be supported by evidence extracted from the original text data, it is considered to better reflect the conversational flow of realistic counselors. Conversely, if most responses are generated predominantly by the model without effectively leveraging the given context, such data entries require further processing. This can be achieved by using an additional prompt to enhance the integration of factual evidence and incorporate content-relevant information from the original context without compromising fluency. Third, to enhance the quality of the multi-turn dialogue data further, we utilize a prompt to revise the data in terms of empathy, supportiveness, guidance, and safety. Ultimately, we perform manual proofreading to further ensure data quality and to derive the processed multi-turn dialogue data.
A example is shown in Figure 4, with three steps to generate high-quality data. The details of the prompt used in this part is shown in the Appendix A.1. In fact, we have compared using a single complex prompt to achieve the same goal as the aforementioned three steps. However, the utilization of a single prompt often results in an abundance of model-generated templated responses, rather than the comprehensive integration of provided factual information. This is the primary reason we opt for a pipeline approach to generate data. The statistics of the data is reported in Table 1.
@@ 2.2 Multi-Turn Dialogue Construction
2.3 knowledge-based QA
In addition to the daily QA interactions, we also introduce some more abstract psychological knowledge data, such as explanations of psychological terms. We crawl books related to psychology from the web and then use Qwen-72B to extract knowledge-based QA from them. Specifically, we segment books into text spans using a predefined fixed length, identifying the nearest sentence or paragraph as segmentation indicators. These text spans serve as the fundamental units for subsequent QA generation through the utilization of LLMs. First, the LLM generates questions and their corresponding answers. These question-answer pairs are then input into two LLM-based student modules, one utilizing retrieval-augmented generation (RAG) and the other without RAG, to produce two new sets of answers. Subsequently, a teacher module, also based on an LLM, evaluates and selects best answer from those generated by the student modules. Furthermore, to ensure the quality and accuracy of the generated QA pairs, a manual validation process is implemented, wherein human evaluators assess and eliminate low-quality data. In addition, we extract after-school exercises from several books and convert them to QA format, with the corresponding comprehensive answer analyses. Finally, we obtain 10K knowledge QA data.
2.4 Supervised Fine-Tuning
After obtaining the high-quality psychological dataset, we further conduct supervised fine-tuning (SFT) to enhance the capability of the LLM in psychology. We select Qwen1.5-14B-Chat as our backbone model, a decoder-only architecture demonstrating excellent performance in both English and Chinese. We then perform SFT on the top of Qwen1.5-14B-Chat to develop PsycoLLM. We employ cross-entropy as the objective function to maximize the likelihood probability. The hyperparameters in this process are configured as follows: the learning rate is set to 1e-5, the batch size is set to 128, and the number of epochs is set to 3. We employ full parameter fine-tuning to train the model using eight A6000 GPUs, each equipped with 48GB memory.
3 Benchmark
@@
3.1 Dataset Collection and Preprocessing

### v2

@@ Abstract
Mental health has attracted substantial attention in recent years and LLM can be an effective technology for alleviating this problem owing to its capability in text understanding and dialogue. However, existing research in this domain often suffers from limitations, such as training on datasets lacking crucial prior knowledge and evidence, and the absence of comprehensive evaluation methods. In this paper, we propose a specialized psychological large language model (LLM), named PsycoLLM, trained on a proposed high-quality psychological dataset, including single-turn QA, multi-turn dialogues and knowledge-based QA. Specifically, we construct multi-turn dialogues through a three-step pipeline comprising generation, evidence judgment, and refinement. We augment this process with real-world psychological case backgrounds extracted from online platforms, enhancing the relevance and applicability of the generated data. Additionally, to compare the performance of PsycoLLM with other LLMs, we develop a comprehensive psychological benchmark based on authoritative psychological counseling examinations in China, which includes assessments of professional ethics, theoretical proficiency, and case analysis. The experimental results on the benchmark illustrates the effectiveness of PsycoLLM, which demonstrates superior performance compared to other LLMs.
Index Terms:
Large language model, psychological understanding, psychological evaluation, mental healthI Introduction
@@
Large language models (LLM) have attracted tremendous attention and experienced promising advancements, such as ChatGPT111https://chat.openai.com/, KimiChat222https://kimi.moonshot.cn/, PaLM [1] , BaiChuan [2] and LLama [3]. These models have demonstrated significant advancements across a variety of tasks, including question answering (QA), summarization [4], and information extraction [5]. Moreover, large language models (LLMs) have been effectively applied to diverse domains, such as the medical and legal domains [6, 7, 8]. In recent years, the escalating pressures from personal, and occupational demands have emerged as significant challenges to the mental health status of individuals. Since LLMs are trained on massive data and contain huge number of parameters, many experiments have shown that these models have the strong ability to understand the natural language and provide resaonable feedback. Hence, leveraging LLM to aid individuals in addressing psychological issues may represent an efficacious strategy for mitigating the impact of adverse mental health conditions and making it easier for seekers to obtain timely help.
@@
Many studies have been proposed in this area. For example, The work [9] combined pre-trained LLMs with real-world professional QA data, with the purpose of providing online consultation services. [10] and [11] constructed a multi-turn empathetic conversation dataset and used a base LLM for fine-tuning. Further, EmoLLM [12] follow a similar way and fine-tune the open base model on the mental health dataset. Although these existing efforts have contributed significantly to the advancement of the field, they still exhibit certain limitations. First, most existing psychological fine-tuning datasets are generated by even larger language model such as ChatGPT. However, the generation process lacks adequate prior knowledge and evidence, potentially leading to the creation of datasets that deviate from real-world psychological communication scenarios. Second, owing to the lack of a comprehensive psychological benchmark, the evaluation of most studies in this area often relies on expert assessment based on subjective judgment or overly depends on other large models, which may not fully encompass the subtle performance. For instance, PsyBench [13] utilized questions generated by GPT-4 to evaluate model performance.
@@
Therefore, in this paper, we propose a psychological LLM (PsycoLLM) trained on a high-quality dataset tailored to the domain of psychology, encompassing single-turn QA, multi-turn dialogues, and knowledge-based QA. Specifically, for single-turn QA, we collect question-answer pairs from open platforms. For multi-turn dialogues, we develop an innovative pipeline with generation, evidence judgment, and refinement for producing high-quality data. For knowledge-based QA, we extract after-school exercises from psychology textbooks and use LLMs enhanced with agents to extract QA pairs from psychology books. Additionally, to explore the performance of the proposed model and exiting LLMs, we introduce a benchmark based on authoritative psychological counseling examinations in China for comparative analysis. The benchmark consists of three distinct components: professional ethics, theoretical proficiency, and case analysis. Professional ethics and theoretical proficiency are assessed through multiple-choice questions (MCQs) to evaluate the mastery of psychological principles, while the case analysis component emphasizes practical utility by incorporating both MCQs and QA sections to explore the application of the model in real-world scenarios.
The results reveal several insightful observations: I) Some LLMs have achieved an accuracy rate surpassing 60% in MCQs, showing the potential of using LLM to provide psychological assistance, with the proposed PsycoLLM demonstrating superior performance among these models. II) LLMs display greater proficiency in professional ethics compared to their mastery of psychological theory.
@@
II Related Work
II-A Large Language Model
Owing to highly parallelizable architecture of self-attention mechanisms, Transformer-based models have been popular in recent years. BERT [14] was a first pre-trained Transformer-based language model, which trained a bidirectional language encoder with specially designed masked language modelling. GPT [15] trained a Transformer-based decoder by generative pre-training. Subsequently, GPT-2 [16] was introduced as an expanded version of GPT, featuring a similar architecture but with a significantly larger scale, encompassing up to 1.5 billion parameters. BART [17] used a sequence-to-sequence architecture and combined the strengths of BERT and GPT, which employed a bidirectional encoder similar to BERT and an autoregressive decoder similar to GPT. These models typically followed a pre-training and fine-tuning paradigm, which need to fine-tune these model for different downstream tasks. Owing the findings that the larger models typically yield better performance, LLM become a prominent topic of research and have achieved significant success. For example, GPT-3 [18] with 175B parameters surpassed its predecessor, GPT-2, and performed well in few-shot learning. Recently, ChatGPT based on GPT-3.5, trained with a sizable amount of internet-sourced text data, has drawn sustainable attention which is exceptional at producing human-like responses. Building upon this success, OpenAI introduced GPT-4 [19], which further enhanced the capabilities of its predecessors. In contrast to ChatGPT and GPT-4 were proprietary models, LLaMA [3] was a open-resource language model, which attracted widespread attention from both academia and industry due to its accessible nature and has achieved noteworthy performance. Further, Touvron et al [20] proposed an updated version of LLamA, called LLaMA2, which was trained on a larger corpus and outperformed LLaMA1 in many benchmarks. Additionally, many other LLMs based on LLaMA architecture also have been proposed, including Alpaca [21] and Mistral-7B [22], where the former first employed instruct-following to train the LLaMA and the latter is a 7B-parameter and outperformed LLamA-2-13B in various evaluations. Baichuan [2] introduced a series of large-scale multilingual language models containing 7 billion and 13 billion parameters, trained from scratch on a corpus of 2.6 trillion tokens. ChatGLM [23] was an evolving family of large language models, which included ChatGLM [24, 25], ChatGLM2 and ChatGLM3, repectively. Besides, the Yi series models [26] were also trained from scratch on 3T multilingual corpus and notable for their ability to extend context length up to 200K by continuing pretraining. Furthermore, Qwen [27] represented another open-resource LLMs and obtained superior performance across a multitude of downstream tasks, especially in Chinese, which was also the main reason we choose it as our backbone.
II-B Benchmark for Mental Health Support
Many studies have been devoted to evaluating the performance of the model across various domains. For example, in the general domain, [28] proposed a Chinese evaluation benchmark to evaluate advanced knowledge and reasoning abilities of Chinese-based models. Zhong at al [29] designed a human-centric benchmark for assessing the performance of foundation models via 20 diverse tasks across a wide variety of subjects. The MMLU benchmark [30] provided a comprehensive evaluation benchmark encompassing various tasks to evaluate the performance of a text model in multi-task contexts. In the medical domain, [31] proposed a localized medical benchmark across various clinical medical professions to evaluate the existing LLMs for medicine. In the legal domain, [32] constructed a Chinese legal benchmark based on the logic of legal practice, categorizing the legal capabilities of language models into different levels. In the psychological domain, several studies have introduced benchmarks to assess model performance in mental health contexts. Jin et al [33] introduced a meticulously crafted benchmark to evaluate the model performance in five mental health-related sub-tasks. Zhang et al [13] used GPT-4 to generate question and constructed a psychological benchmark aimed at evaluating model performance in psychology. Furthermore, Qiu et al [34] created a benchmark with fine-grained labels for each dialogue and focused on dialogue safety in mental health support. Despite the advances made with these benchmarks, there are several limitations. Some benchmarks narrow their scope to specific subfields or isolated tasks within mental health, neglecting to evaluate the models’ overall psychological knowledge mastery. Furthermore, benchmarks generated by large models carry the risk of deviating from real-world contexts. Compared to these works, we propose a benchmark that incorporates different levels of official examinations in China, enabling a comprehensive assessment of both theoretical knowledge and practical proficiency.
II-C Mental Application in LLM
The success of LLMs in general domains has sparked interest in their application across various vertical fields. Many LLMs tailored for specific domains have achieved remarkable success [35, 36]. For instance, HuatuoGPT [37], an LLM designed for the medical domain, has demonstrated excellent performance in medical consultations. Based on HuatuoGPT, HuatuoGPT-II [38] introduced a unified domain adaptation protocol to combine continued pre-training and supervised fine-tuning into a single process, which achieved better performance than HuatuoGPT. In the legal domain, Yue et al [39] developed an LLM tailored for building intelligent legal systems with legal reasoning and knowledge retrieval capability. Mental health is another crucial domain that has garnered significant attention in many studies [40, 41, 42, 43, 44, 45, 46]. For example, the work [47] developed an LLM-based expert-guided pipeline to analyze raw interview transcripts and extract insights. Zhu et al [48] investigated the use of LLM to perform inferring users’ underlying goals and fundamental psychological needs, finding that LLM-based approaches can be comparable to human performance. Wu et al [49] introduced an anthropomorphic LLM-based conversational agent designed to offer personalized support and suggest practical actions based on positive psychology and social psychology research.
Several LLMs have also been proposed to aid help-seekers in accessing effective mental health support. Lai et al [9] proposed an AI-based assistive tool that utilized LLMs for QA, aiming to alleviate the demand on mental health professionals. PsyQA [50] introduced a high-quality Chinese dataset of psychological health support in the form of question and answer pair. Based on PsyQA, Qiu et al [11] further used ChatGPT to covert single-turn dialogues to multi-turn dialogues and subsequently fine-tuned ChatGLM2-6B on this enhanced dataset. SoulChat [10] constructed a multi-turn empathetic conversation dataset and fine-tuned ChatGLM with this data. Yang et al [44] formalized the interpretable mental health analysis task and built a multi-task and multi-source interpretable mental health instruction dataset. After that, based on the contructed dataset, they released MentaLLaMA, an open-source instruction-following LLM fine-tuned on LLaMA2. The work [51] fine-tuned the LLM across multiple datasets simultaneously, demonstrating its effectiveness in various mental health prediction tasks. Although these efforts have greatly advanced the field, they rely too heavily on distillations from other models during data generation, leading to the creation of data that may not accurately reflect real-world scenarios. In this paper, we constructed a high-quality psychological dataset, including single-turn, multi-turn dialogue data, and knowledge QA data.
III PsycoLLM
@@
In this section, we present the details of dataset including, single-turn QA, multi-turn dialogue and knowledge-based QA, respectively. The overview of dataset are shown in Fig. 1. Besides, we give the training process of PsycoLLM.
III-A Single-Turn QA Construction
@@
There are several publicly accessible websites committed to establishing a psychology platform and offering online solutions for individuals seeking psychological assistance, such as Yixinli, Zhihu and so on. The data from these platforms can be regarded as real-world inquiries and the solutions provided by professionals in response to these inquiries. On the original website, users typically provide descriptions along with their questions, including background information. Subsequently, professional or experienced individuals offer detailed solutions or advice in response to these questions. We collect over 267,000 pairs of data from websites. Furthermore, several important data cleaning procedures are implemented to enhance the quality and relevance of the dataset, including:
Removal of irrelevant content: This involves eliminating extraneous materials such as advertisements, which do not contribute to the core data and could potentially introduce noise.
Deletion of short data entries: Data entries with fewer than 100 characters are removed to ensure that the dataset contains sufficiently substantive content for meaningful analysis.
Exclusion of low-engagement answers: Answers with fewer than 5 likes are excluded to prioritize content that has been deemed valuable by users. The number of likes functions as a measure of public endorsement of the response. A high number of likes suggests a broad consensus among users regarding the validity and reliability of the answer.
Exclusion of responses from lower-level counselors or individual answers: This ensures the dataset reflects high-quality professional advice to some extent. The respondents’ level may serve as an indicator of their extensive experience in assisting help-seekers, which, in turn, can be indicative of their professional competence. This metric provides a basis for assessing the quality of the provided information.
@@
After these data cleaning procedures, we obtain over 155k pairs for single-turn QA. We report detailed statistics of the processed data drawn from different aspects. We divide the processed data into 9 major topics and 127 subtopics, as illustrated by the distribution of major topics depicted in Fig. 2. It is observed that the number of reports related to emotional issues and regulation, as well as interpersonal relationships and socializing, is higher than those in other categories, each comprising more than 20% of the total. Following these topics, family and marriage, and personal growth and development are the next most significant areas, each accounting for more than 10% of the reports. Besides, the visualization of the collected data is depicted through word clouds, as shown in Fig. 3, which provides an intuitive depiction of the frequency and prominence of specific terms within the dataset. The analysis underscores the pervasive connection of the majority of these data to everyday psychological counseling, further emphasizing the relevance and reasonability of the collected data.
| Category | Prompt Template | |||||||
| Multi-Turn Generation |
| Evidence Support |
| Dialogue Evaluation |
@@
III-B Multi-Turn Dialogue Construction
Based on the extracted data in section III-A, we select a subset of the highest upvoted answers for each question to generate multi-turn dialogue data using KimiChat. This process aims to emulate the behavior of a psychological counselor who progressively asks probing questions to gain insight into the user’s psychological state and requirements, thereby facilitating more effective assistance. The multi-turn dialogue mainly includes two roles: the person seeking help and psychological professionals. We employ a three-step pipeline to generate high-quality multi-turn dialogue data. First, we use an appropriate prompt to guide the KimiChat to construct multi-turn dialogue data between pre-defined roles according to the selected pairs. The original pairs serve as prior knowledge to guide KimiChat in generating multi-turn conversations, thereby enhancing the fidelity of the generated data to real-world conversations. Second, we employ an additional prompt to assess whether the answers in the multi-turn dialogue dataset are derived from the original context. If the majority of responses in the multi-turn dialogue can be supported by evidence extracted from the original text data, it is considered to better reflect the conversational flow of realistic counselors. Conversely, if most responses are generated predominantly by the model without effectively leveraging the given context, such data entries require further processing. This can be achieved by using an additional prompt to enhance the integration of factual evidence and incorporate content-relevant information from the original context without compromising fluency. Third, to enhance the quality of the multi-turn dialogue data further, we utilize a prompt to revise the data in terms of empathy, supportiveness, guidance, and safety. Ultimately, we perform manual proofreading to further ensure data quality and to derive the processed multi-turn dialogue data. The prompts used for multi-turn dialogue generation, evidence support identification, and dialogue evaluation are presented in Table I.
A example is shown in Fig. 4, with three steps to generate high-quality data. In fact, we have compared using a single complex prompt to achieve the same goal as the aforementioned three steps. However, the utilization of a single prompt often results in an abundance of model-generated templated responses, rather than the comprehensive integration of provided factual information. This is the primary reason we opt for a pipeline approach to generate data. The statistics of the data is reported in Table II.
III-C knowledge-based QA
In addition to the daily QA interactions, we also introduce some more abstract psychological knowledge data, such as explanations of psychological terms. We crawl books related to psychology from the web and then use Qwen-72B to extract knowledge-based QA from them. Specifically, we segment books into text spans using a predefined fixed length, identifying the nearest sentence or paragraph as segmentation indicators. These text spans serve as the fundamental units for subsequent QA generation through the utilization of LLMs. First, the LLM generates questions and their corresponding answers. These question-answer pairs are then input into two LLM-based student modules, one utilizing retrieval-augmented generation (RAG) and the other without RAG, to produce two new sets of answers. Subsequently, a teacher module, also based on an LLM, evaluates and selects best answer from those generated by the student modules. Furthermore, to ensure the quality and accuracy of the generated QA pairs, a manual validation process is implemented, wherein human evaluators assess and eliminate low-quality data. In addition, we extract after-school exercises from several books and convert them to QA format, with the corresponding answer analyses. Finally, we obtain 10K knowledge QA data.
III-D Supervised Fine-Tuning
After obtaining the high-quality psychological dataset, we further conduct supervised fine-tuning (SFT) to enhance the capability of the LLM in psychology. We select Qwen1.5-14B-Chat as our backbone model, a decoder-only architecture demonstrating excellent performance in both English and Chinese. We then perform SFT on the top of Qwen1.5-14B-Chat to develop PsycoLLM.
At training stage, assuming the text input as a sequence, e.g., , where each is a text token and is the length of . The core architecture of Qwen is based on Transformer-Decoder structure [52], which is a typical autoregressive framework designed to sequentially predict each subsequent word in a given sequence. This can be mathematically formulated as follows:
For training PsycoLLM, we utilize the cross-entropy loss function as the objective. The model is trained to maximize the negative conditional log-likelihood of the predicted sequence given the input data.
where is the trainable parameters of the model.
The hyperparameters in this process are configured as follows: the learning rate is set to 1e-5, the batch size is set to 128, and the number of epochs is set to 3. We employ full parameter fine-tuning to train the model using eight A6000 GPUs, each equipped with 48GB memory.
IV Benchmark
IV-A Dataset Collection and Preprocessing